nx+12=3y what is n if the slope is 4
Answer by checkley71(8403)

NX+12=3Y
3Y=NX+12
Y=NX/3+12
FOR THE SLOPE TO BE = TO 4 THEN N/3=4 OR N=12.
Y=12X/3+12
Y=4X+12